OpenStudy (anonymous):

multiply and simplify sqrt6(sqrt3+5sqrt2)

OpenStudy (***[isuru]***):

\[\sqrt{6} ( \sqrt{3} + 5\sqrt{2}\]\[(\sqrt{6} \times \sqrt{3} )+ (\sqrt{6} \times 5\sqrt{2})\]\[\sqrt{18} + 5\sqrt{12}\]ALSO..\[\sqrt{18 } = \sqrt{9 \times 2} =3 \sqrt{2}\]\[5\sqrt{12} = 5\sqrt{4 \times 3} = 12\sqrt{3}\] which means \[3\sqrt{2} + 10\sqrt{3}\] hope this will help ya!!